# Displacement-Squeeze receiver for BPSK displaced squeezed vacuum states surpassing the coherent-states Helstrom bound under imperfect conditions

### v1

@@ IV.2 Phase Diffuse Noise in communication channel
Fig. 8(c) and (b) depict the performance of DSR as a function of the mean signal photon number under a phase diffusion strength of $\sigma=0.1$. The results show that DSR can surpass the SQL of the S-BPSK scheme within a certain photon-number range ($N\in\left(0.28,0.93\right)$), achieving a maximum advantage of 3.32 dB beyond the SQL. However, at high photon numbers, DSR no longer exceeds the SQL. In particular, when the optimal threshold reaches the maximum resolution limit of the PNR detector, the error probability does not saturate but instead increases with the mean photon number. Furthermore, we observe that a PNR detector with a resolution of 2n (even) provides only marginal improvement in DSR performance compared to one with a resolution of 2n-1; no pronounced oscillatory drop in error probability is evident. This behavior is consistent with the variation in the optimal threshold, as shown in panel (c).
@@ IV.3 Thermal noise in the receiver
is the circularly symmetric complex Gaussian distribution. Here $n_{t}$ is the mean photon number of the thermal background mode, which under thermal equilibrium is
$n_{t}=\left[\exp\!\left(\frac{h\nu}{kT_{0}}\right)-1\right]^{-1},$ (35)
with $h$ the Planck constant, $k$ the Boltzmann constant, $\nu$ the optical frequency, and $T_{0}$ the receiver temperature. In practical fiber-optic communication systems, the receiver-originated background often dominates over the channel-induced thermal noise [9].
Consistent with [31, 29], we assume that the thermal-noise contamination occurs immediately before the PNR detector, as illustrated in Fig. 9.
Figure 9: Schematic of the DSR under receiver thermal noise. The processed states $\zeta_{i}$ impinging on the PNR(M) are contaminated by receiver thermal noise and become $\zeta_{i}^{\prime}$.
Remark (coherent state $\Rightarrow$ displaced thermal state): Under the channel $\Phi_{n_{t}}$ in Eq. (33), a coherent-state input becomes a displaced thermal state. Specifically, using the Weyl relation $D(\lambda)D(\alpha)=e^{(\lambda\alpha^{\ast}-\lambda^{\ast}\alpha)/2}D(\lambda+\alpha)$, one verifies the displacement covariance
@@
Consequently, the states $\zeta_{i}$ measured by PNR(M), originally $\left|0\right\rangle$ and $\left|2\alpha\mathrm{e}^{r}\right\rangle$, are transformed into
@@
$\zeta_{1}^{\prime}$ $=\Phi_{n_{t}}\!\big(|2\alpha\mathrm{e}^{r}\rangle\langle 2\alpha\mathrm{e}^{r}|\big)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,\left|\lambda+2\alpha\mathrm{e}^{r}\right\rangle\left\langle\lambda+2\alpha\mathrm{e}^{r}\right|.$
Equivalently, by Eq. (37), one may write $\zeta_{0}^{\prime}=\rho_{\mathrm{th}}(n_{t})$ and $\zeta_{1}^{\prime}=D(2\alpha\text{e}^{r})\rho_{\mathrm{th}}(n_{t})D^{\dagger}(2\alpha\text{e}^{r})$.
@@
$q_{n}^{M}(\zeta_{i}^{\prime})=\mathrm{Tr}\!\left(\Pi_{n}\,\zeta_{i}^{\prime}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,q_{n}^{M}[\mu_{i}(\lambda)],$ (39)
where $\Pi_{n}$ is the POVM element corresponding to outcome $n$, and $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\alpha\mathrm{e}^{r}|^{2}$. Substituting Eq. (39) into Eq. (17) and Eq. (19) yields the optimal threshold $n_{\text{th}}^{*}$ and the error probability $P_{\text{err}}^{\text{DSR}}$ of the DSR in the presence of receiver thermal noise, respectively.
@@
Figure 10: Simulation results of DSR with PNR(M) under thermal noise condition: (a) Error probability ($n_{t}=0.001$), (b) Error probability ($M=10$), (c) Ratio (in dB) of the benchmark $P_{\text{err}}^{\text{DSS}}$ ($M=10$), (d) Optimal threshold ($M=10$), (e) Fock-basis population of $\zeta_{i}^{\prime}$ (the input energy $N=1.0$ and $n_{t}=0.001$), and (f) Maximum tolerable thermal noise $n_{t}^{\text{max}}$ for different photon number resolution M.
@@
Fig. 10(a) and (b) depict the error probability of the DSR as a function of the mean signal photon number, under different PNR detector resolutions and various thermal noise strengths. The error probability curves exhibit an oscillatory pattern and decrease overall with increasing signal energy, which is a consequence of the continuous adjustment of the optimal threshold $n_{\text{th}}^{*}$ (see panel (d)). This trend continues until $n_{\text{th}}^{*}$ reaches the maximum resolution of the PNR detector, after which the error probability saturates and no longer varies with the mean photon number. Fig. 10(c) shows the ratio (in dB) of the DSR error probability to the SQL of standard BPSK versus the mean signal photon number under different noise strengths. It can be observed that the DSR can still surpass the SQL at certain noise levels. For instance, when $n_{t}=10^{-3}$, the error probability of the DSR with a PNR(M) detector is lower than the SQL by -6.34dB (M=1), -13.43dB (M=2), -20.79dB (M=3), and -28.27dB (M=4).
Furthermore, we note that $n_{\text{th}}^{*}$ increases in steps of one until it attains the maximum resolution of the PNR detector. This occurs because the Fock-basis population of the thermal-noise-perturbed state $\zeta_{0}^{\prime}$ still decreases monotonically with increasing photon number. This behavior differs from that under phase-diffusion noise (see Fig. 8(d)).
@@
We then introduce the maximum tolerable thermal noise $n_{t}^{\text{max}}$ as a figure of merit, defined as the highest noise level for which $P_{\text{err}}^{\text{DSR}}\leq P_{\text{SQL}}^{\text{DSS}}$ at a given signal energy N. This quantity is shown in Fig. 10(e). Thus, the DSR outperforms the SQL when $n_{t}<n_{t}^{\text{max}}$, corresponding to the region below the $n_{t}^{\text{max}}$ curve. The plot displays M distinct peaks, after which it decays toward zero.
@@ V Conclusion
We presented a displacement-squeeze receiver for S-BPSK discrimination based on Gaussian preprocessing and PNR/MAP thresholding. In the ideal, equal-prior case, $n_{\text{th}}^{*}=1$ and $P_{\text{err}}^{\text{DSR}}=\frac{1}{2}\exp\left[-4N\left(N+1\right)\right]$, independent of the PNR number resolution. The resulting performance is provably within a constant factor of the S-BPSK Helstrom bound for all N, $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$, while surpassing the S-BPSK SQL at $N\approx 0.3$ and the C-BSPK Helstrom bound at $N\approx 0.4$ (sub-1% errors near $N\approx 0.6$). Under nonideal operation, we modeled detector inefficiency and dark counts, channel phase diffusion, and receiver thermal noise, and found that adaptive MAP thresholds maintain an advantage over the SQL within practical ranges (e.g., under $\sigma=0.1$, DSR exceeds the S-BPSK SQL for $N\in(0.28,0.93)$).
Future work includes extending the architecture to multi-ary constellations (e.g., QPSK and 16QAM) and validating performance in hardware under dynamic channel conditions.

### v2

@@ IV.2 Phase Diffusion Noise in communication channel
However, at high photon numbers, IS-Kennedy no longer exceeds the SQL. In particular, when the optimal threshold reaches the maximum resolution limit of the PNR detector, the error probability does not saturate but instead increases with the mean photon number. Furthermore, we observe that a PNR detector with a resolution of 2n (even) provides only marginal improvement in IS-Kennedy performance compared to one with a resolution of 2n-1 (odd); no pronounced oscillatory drop in error probability is evident. This behavior is consistent with the variation in the optimal threshold, as shown in panel (c).
@@ IV.3 Thermal noise in the receiver
is the circularly symmetric complex Gaussian distribution. Here $n_{t}$ is the mean photon number of the thermal background mode, which under thermal equilibrium is $n_{t}=\left[\exp\!\left(\frac{h\nu}{kT_{0}}\right)-1\right]^{-1}$ with $h$ the Planck constant, $k$ the Boltzmann constant, $\nu$ the optical frequency, and $T_{0}$ the receiver temperature. In practical fiber-optic communication systems, the receiver-originated background often dominates over the channel-induced thermal noise [16].
Figure 9: Schematic of the IS-Kennedy under receiver thermal noise. The processed states $\zeta_{i}$ impinging on the PNR(M) are contaminated by receiver thermal noise and become $\zeta_{i}^{\prime}$.
Consistent with Ref. [41, 40], we assume that the thermal-noise contamination occurs immediately before the PNR detector, as illustrated in Fig. 9.
Note (coherent state $\Rightarrow$ displaced thermal state): Under the channel $\Phi_{n_{t}}$ in Eq. (38), a coherent-state input becomes a displaced thermal state. Specifically, using the Weyl relation $D(\lambda)D(\alpha)=e^{(\lambda\alpha^{\ast}-\lambda^{\ast}\alpha)/2}D(\lambda+\alpha)$, one verifies the displacement covariance
@@
[…]
@@
[…]
Equivalently, by Eq. (41), one may write $\zeta_{0}^{\prime}=\rho_{\mathrm{th}}(n_{t})$ and $\zeta_{1}^{\prime}=D(2\gamma)\rho_{\mathrm{th}}(n_{t})D^{\dagger}(2\gamma)$.
@@
$P\left(n|i\right)=\text{Tr}\!\left(\Pi_{n}^{\text{ideal}}\,\zeta_{i}^{\prime}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\cdot f_{n_{t}}(\lambda)\cdot\text{Poiss}\left\{n;\mu_{i}(\lambda)\right\}$ (43)
where $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\gamma|^{2}$. Substituting Eq. (43) into the MAP rule in Eq. (16), we numerically obtain the optimal decision rule and the corresponding error probability in the presence of receiver thermal noise. A general closed-form monotonicity proof for the likelihood ratio is not available in this mixed-state setting; accordingly, the threshold values reported below are numerically verified MAP-equivalent thresholds within the explored parameter range.
@@
Figure 10: Simulation results of IS-Kennedy with PNR(M) under thermal noise condition: (a) Error probability ($n_{t}=0.001$), (b) Error probability ($M=10$), (c) Ratio (in dB) of the benchmark $P_{\text{SQL}}^{\text{S-BPSK}}$ ($M=10$), (d) Optimal threshold ($M=10$), (e) Fock-basis population of $\zeta_{i}^{\prime}$ (the input energy $N=1.0$ and $n_{t}=0.001$), and (f) Maximum tolerable thermal noise $n_{t}^{\text{max}}$ for different photon number resolution M.
@@
Fig. 10(a) and (b) depict the error probability of the IS-Kennedy as a function of the mean signal photon number, under different PNR detector resolutions and various thermal noise strengths. The error probability curves exhibit an oscillatory pattern and decrease overall with increasing signal energy, which is a consequence of the continuous adjustment of the optimal threshold $n_{\text{th}}^{*}$ (see panel (d)). This trend continues until $n_{\text{th}}^{*}$ reaches the maximum resolution M of the PNR detector, after which the error probability saturates and no longer varies with the mean photon number. As analyzed in Sec. IV.A, the saturated error probability is dominated by the false alarm probability $P_{\text{FA}}$ (Eq. (20)), and when the signal energy tends to infinity, the optimal threshold must be equal to M. Therefore, the saturated error probability is:
[…]
$=\frac{1}{2}\sum_{n=\text{M}}^{\infty}\frac{{n_{t}}^{n}}{(1+n_{t})^{n+1}}=\frac{1}{2}\left(\frac{n_{t}}{1+n_{t}}\right)^{\text{M}}$
Fig. 10(c) shows the ratio (in dB) of the IS-Kennedy error probability to the SQL of standard BPSK versus the mean signal photon number under different noise strengths. It can be observed that the IS-Kennedy can still surpass the SQL at certain noise levels. For instance, when $n_{t}=10^{-3}$, the error probability of the IS-Kennedy with a PNR(M) detector is lower than the SQL by -6.34dB (M=1), -13.43dB (M=2), -20.79dB (M=3), and -28.27dB (M=4).
@@
We then introduce the maximum tolerable thermal noise $n_{t}^{\text{max}}$ as a figure of merit, defined as the highest noise level for which $P_{\text{err}}^{\text{ISK}}\leq P_{\text{SQL}}^{\text{S-BPSK}}$ at a given signal energy N. This quantity is shown in Fig. 10(f). Thus, the IS-Kennedy outperforms the SQL when $n_{t}<n_{t}^{\text{max}}$, corresponding to the region below the $n_{t}^{\text{max}}$ curve. The plot displays M distinct peaks, after which it decays toward zero.
IV.4 Transmission loss in the communication channel
A practically important limitation of squeezed-state communication is transmission loss. In contrast to the ideal analysis of Sec. III, propagation through a lossy channel does not merely reduce the signal displacement, but also degrades the squeezing resource itself by mixing the signal mode with an environmental vacuum mode. Equivalently, in the Stinespring representation, a pure-loss bosonic channel can be modeled by a beam splitter of transmittance $T$ that couples the signal to an environmental vacuum state [10, 37].
For nonclassical inputs such as displaced squeezed states, this interaction generally generates system-environment entanglement. However, since the environment is inaccessible to the receiver, the relevant state for discrimination is the reduced output state of the signal mode, obtained after tracing out the environment. Therefore, the received signals are correctly described as single-mode mixed Gaussian state, namely a displaced squeezed thermal state (DSTS), rather than pure displaced squeezed vacuum state [37].
Figure 11: Schematic diagram of pure-loss channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{\text{loss}}$ (mixed state) in the pure-loss communication channel, which $T$ is the equivalent transmittance.
1. Input Gaussian-state description.
[…]
$\vec{d_{i}}=\begin{pmatrix}(-1)^{i+1}\sqrt{2}\alpha\\ 0\end{pmatrix},\qquad\mathbf{V}=\frac{1}{2}\begin{pmatrix}e^{-2r}&0\\ 0&e^{2r}\end{pmatrix}$. (45)
2. Pure-loss channel as a Gaussian map.
We model propagation by a pure-loss bosonic channel with transmittance $T\in[0,1]$ (see Fig. 11). At the operator level, the channel can be represented as
$\hat{a}_{\mathrm{out}}=\sqrt{T}\,\hat{a}_{\mathrm{in}}+\sqrt{1-T}\,\hat{v}$, (46)
where $\hat{v}$ is an environmental vacuum mode [37]. For displaced squeezed vacuum state, this channel acts linearly on first and second moments:
$\vec{d_{i}}^{(T)}=\sqrt{T}\,\vec{d_{i}},\qquad\mathbf{V}^{(T)}=T\cdot\mathbf{V}+(1-T)\frac{\mathbf{I}}{2}=\frac{1}{2}\begin{pmatrix}1-T+Te^{-2r}&0\\ 0&1-T+Te^{2r}\end{pmatrix}$. (47)
Several remarks are immediate: (i) When $T=1$, one recovers the ideal pure displaced squeezed states; (ii) For $0<T<1$ and $r\neq 0$, the determinant of $\mathbf{V}^{(T)}$ is larger than $1/4$, which means that the reduced output state is mixed. Therefore, after propagation, the received S-BPSK signals are no longer pure displaced squeezed vacuum states, but displaced squeezed thermal states. This is precisely the local manifestation of the system–environment coupling discussed above: the information leaked to the inaccessible environment appears at the receiver as a loss-induced mixedness.
3. Equivalent squeezed-thermal parametrization.
It is convenient to rewrite $\mathbf{V}^{(T)}$ in the standard squeezed-thermal form
$\mathbf{V}^{(T)}=\left(n_{T}+\frac{1}{2}\right)\begin{pmatrix}e^{-2r_{T}}&0\\ 0&e^{2r_{T}}\end{pmatrix}$. (48)
By comparing matrix elements, the residual squeezing parameter $r_{T}$ and the equivalent thermal photon number $n_{T}$ are obtained as
$r_{T}$ $=\frac{1}{4}\ln\!\left(\frac{1-T+Te^{2r}}{1-T+Te^{-2r}}\right),$ (49)
$n_{T}$ $=\frac{1}{2}\left[\sqrt{(1-T+Te^{-2r})(1-T+Te^{2r})}-1\right].$
Therefore, the input of IS-Kennedy is:
$\rho_{i}^{\text{loss}}=\left\{\begin{aligned} &D(-\sqrt{T}\alpha)S\left(r_{T}\right)\rho_{\text{th}}\left(n_{T}\right)S^{\dagger}\left(r_{T}\right)D^{\dagger}(-\sqrt{T}\alpha),\quad i=0\\ &D(+\sqrt{T}\alpha)S\left(r_{T}\right)\rho_{\text{th}}\left(n_{T}\right)S^{\dagger}\left(r_{T}\right)D^{\dagger}(+\sqrt{T}\alpha),\quad i=1\end{aligned}\right$. (50)
These expressions have a clear physical meaning. When $T\to 1$, one has $r_{T}\to r$ and $n_{T}\to 0$, so the ideal case is recovered. When $0<T<1$, the effective squeezing is reduced from $r$ to $r_{T}$, while a nonzero thermal component $n_{T}$ appears. When $T\to 0$, the signal is completely extinguished and both hypotheses collapse to vacuum. Hence, transmission loss affects the receiver in two distinct ways: it attenuates the displacement and, at the same time, destroys the purity of the squeezing resource.
4. Matched IS-Kennedy processing after a lossy channel.
In the ideal receiver of Sec. III, the nulling displacement $D(\alpha)$ and the inverse-squeezing S(r)S(-r) are matched to the transmitted state. After a lossy channel, the received hypotheses are centered at $\pm\sqrt{T}\alpha$, rather than at $\pm\alpha$, and exhibit residual squeezing $r_{T}$ rather than $r$. Therefore, the matched version of the IS-Kennedy in the lossy scenario should use
$D(\alpha)\longrightarrow D(\sqrt{T}\alpha),\quad S(-r)\longrightarrow S(-r_{T}).$ (51)
The nulling displacement $D(\sqrt{T}\alpha)$ maps the attenuated BPSK alphabet into a lossy squeezed OOK alphabet:
$\sigma_{i}=\left\{\begin{aligned} &S\left(r_{T}\right)\rho_{\text{th}}\left(n_{T}\right)S^{\dagger}\left(r_{T}\right),\quad i=0\\ &D(2\sqrt{T}\alpha)S\left(r_{T}\right)\rho_{\text{th}}\left(n_{T}\right)S^{\dagger}\left(r_{T}\right)D^{\dagger}(2\sqrt{T}\alpha),\quad i=1\end{aligned}\right$. (52)
Subsequently, applying the matched inverse-squeezing $S(-r_{T})$ removes the residual anisotropy of the covariance matrix. As in the ideal case, the inverse-squeezing converts the remaining squeezing resource into a larger effective coherent-state displacement. Using the standard transformation of displacement operators under squeezing, one obtains
$S(-r_{T})D(2\sqrt{T}\alpha)S(r_{T})=D(2\sqrt{T}\alpha e^{r_{T}}).$ (53)
Defining
$\gamma_{T}=\sqrt{T}\,\alpha e^{r_{T}},$ (54)
the two output hypotheses impinging on the PNR detector become (an OOK pair composed of a thermal state and a displaced thermal state)
$\zeta_{i}=\left\{\begin{aligned} &\rho_{\text{th}}(n_{T}),\quad i=0,\\ &D(2\gamma_{T})\rho_{\text{th}}(n_{T})D^{\dagger}(2\gamma_{T}),\quad i=1.\end{aligned}\right$. (55)
The above formulas show that transmission loss weakens the squeezing advantage through two simultaneous mechanisms. First, the coherent-state separation gain is reduced from $\gamma=\alpha e^{r}$ to $\gamma_{T}=\sqrt{T}\alpha e^{r_{T}}$. Second, the perfectly nulled hypothesis is no longer vacuum. Under ideal conditions, the nulling branch produces $|\zeta_{0}=\ket{0}\bra{0}$. After transmission loss, however, $\zeta_{0}^{(T)}=\rho_{\mathrm{th}}(n_{T})$.
Hence, the conditional photon-number probabilities are
$P(n|i=0)=\int_{\mathbb{C}}\text{d}^{2}\lambda\cdot f_{n_{T}}(\lambda)\cdot\mathrm{Poiss}(n;|\lambda|^{2}),$ (56)
$P(n|i=1)=\int_{\mathbb{C}}\text{d}^{2}\lambda\cdot f_{n_{T}}(\lambda)\cdot\mathrm{Poiss}(n;|\lambda+2\gamma_{T}|^{2}).$
These expressions are identical in structure to the thermal-noise formulas already used in Sec. IV.C, with the substitutions:
$n_{t}\rightarrow n_{T},\quad\gamma\rightarrow\gamma_{T}.$ (57)
Since the lossy case reduces to the same thermal-state decision structure with the substitutions in Eq. (57), the MAP decision rule and the corresponding error probability are evaluated numerically. As in Secs. IV.B and IV.C, we do not claim a general closed-form proof that the likelihood ratio is monotonic in $n$; the thresholds shown below are therefore numerically verified MAP-equivalent thresholds over the explored parameter range.
For practical optical communication, one may parameterize the transmittance as
$T(a,L)=10^{-aL/10},$ (58)
where $a$ is the attenuation coefficient in dB/km and $L$ is the link length in km. Therefore, once $L$ and $a$ are fixed, the MAP (optimal threshold) and the error probability can be computed numerically without introducing any new statistical machinery. To model transmission loss in a representative long-haul fiber scenario, we fix the attenuation coefficient at a=0.2dB/kma=0.2\ $\text${dB/km}, which is a standard rule-of-thumb value around the 1550 nm telecom window [14].
[…]
Figure 12(a) shows that the error probability of the IS-Kennedy increases monotonically with link length. Compared with the lossy Kennedy receiver, the IS-Kennedy curves exhibit a weak oscillatory structure as the input energy increases, whereas the Kennedy curves remain smooth. This difference originates from the threshold adaptation of the IS-Kennedy: as the received signal energy varies, the optimal threshold changes in discrete steps, which produces the visible ripple-like behavior (see Fig. 12(c)). By contrast, the lossy Kennedy receiver retains the standard on-off structure because a coherent-state input remains coherent under pure loss. A second and more important difference is that, after transmission loss, the nulling branch of the IS-Kennedy is no longer vacuum. For coherent-state signaling, the nulled hypothesis ideally remains vacuum after matched displacement, so the false-alarm probability is determined only by detector imperfections. For squeezed-state signaling, however, the lossy channel transforms the input into a displaced squeezed thermal state; after matched receiver processing, the nulled hypothesis becomes a thermal state rather than vacuum. Consequently, the probability of detecting photons under hypothesis $i=0$ is no longer zero, which explains why loss degrades the IS-Kennedy more strongly than the Kennedy benchmark.
The comparison with the lossy Kennedy receiver in Fig. 12(b) makes this trend more explicit. For low-loss links, the IS-Kennedy retains a clear negative dB ratio over a broad range of $N$, showing that the residual squeezing can still be converted into a receiver-side discrimination gain. As the link length increases, however, the ratio curve shifts upward and eventually approaches or exceeds 0 dB, indicating that the advantage over the Kennedy benchmark is progressively erased by channel attenuation. Therefore, within the fixed source parametrization adopted here, the main benefit of the IS-Kennedy is concentrated in the low-loss regime.
The role of the PNR resolution is illustrated in Fig. 12(d). As analyzed in Sec. IV.C, when the signal energy increases, the optimal threshold increases to M, the miss detection probability $P_{\text{Mi}}$ decreases approaching 0, and the receiver’s error probability is dominated by false alarm probability $P_{\text{FA}}$, so that at very high input energy:
$P_{\text{err}}^{\text{ISK,loss}}\to\frac{1}{2}P_{\text{FA}}(M)=\frac{1}{2}\left(\frac{n_{T}}{1+n_{T}}\right)^{M}$ (59)
However, the difference is that $n_{T}$ varies with the signal energy $N$, so under transmission loss conditions, a saturated error probability does not occur.
Figure 13: Maximum link length $L_{\text{max}}$ of the benchmark $P_{\text{err}}^{\text{K,loss}}$ for different mean photon number $N$ and resolution M
Figure 13 summarizes the practical operating regime through the maximum advantageous link length $L_{\text{max}}$, defined as the longest link for which the IS-Kennedy still outperforms the lossy Kennedy receiver. Hence, the IS-Kennedy is advantageous in the region $L<L_{\text{max}}$, i.e., below the corresponding $L_{\text{max}}$ curve.
Overall, the transmission-loss analysis shows that, within the fixed-$\beta$ S-BPSK parametrization considered here, the IS-Kennedy advantage is concentrated in the low-loss regime. Since the fixed-$\beta=\frac{N}{2N+1}$ might not remain optimal in a lossy channel, this conclusion should not be interpreted as a general exclusion of long-distance links employing squeezed states. We anticipate that allowing $\beta$ to adapt with propagation distance would lower the IS-Kennedy error probability and increase the maximum transmission distance $L_{\text{max}}$, which will be pursued in future work.
@@ V Conclusion
We have proposed and analyzed the inverse-squeezing Kennedy receiver for S-BPSK discrimination. By combining displacement, orthogonally oriented inverse-squeezing, and PNR/MAP detection, the receiver converts transmitter-side squeezing into an effective separation gain at the measurement stage. Under ideal conditions and equal priors, the resulting error probability stays within a constant factor of the S-BPSK Helstrom bound for all N, while surpassing the S-BPSK SQL at $N\approx 0.3$ and the C-BPSK Helstrom bound at $N\approx 0.4$.
Under nonideal conditions, adaptive thresholding mitigates the impact of detector inefficiency, dark counts, phase diffusion, and receiver thermal noise over practical parameter ranges. Transmission loss is more restrictive: it attenuates the signal and degrades the squeezing resource, thereby reducing the advantage over Kennedy-type and coherent-state benchmarks with increasing propagation distance. Therefore, under the fixed launch-state optimization adopted here, the proposed receiver is most favorable in low-loss channels, while its advantage is progressively reduced as propagation loss increases.
Future work should extend the present analysis beyond the fixed-$\beta$ source parametrization adopted here and consider joint optimization of the launch state and receiver settings in the presence of transmission loss. In particular, re-optimizing the squeezing fraction and nulling displacement as functions of channel attenuation may further improve performance and enlarge the advantageous operating regime of the inverse-squeezing Kennedy receiver.